# A Proxy Attack-Free Strategy for Practically Improving the Poisoning Efficiency in Backdoor Attacks

### v2

@@ A-D Attack Success Rate Under Additional Target Classes on the CIFAR-10 Dataset
A-E Experiments on CIFAR-100 Dataset
The results for CIFAR-100 in Table XI are similar to those for CIFAR-10. Our PFS approach outperforms random selection in all cases, with average boosts of 0.047, 0.068, and 0.037 for BadNets, Blended, and Optimized triggers, respectively. Compared to FUS, PFS achieves better attack results in 49 out of 54 settings.
TABLE XI: The attack success rate on CIFAR-100 dataset. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets, Blended, and Optimized are $1.5\%$, $1.5\%$, and $0.75\%$, respectively. Among four select strategies, the best result is denoted in boldface while the underline indicates the second-best result. The settings corresponding to the gray background are represented as the proxy poisoning attack within the attacker phase and the actual poisoning process within the victim phase being consistent in FUS.
@@ A-E Experiments on CIFAR-100 Dataset
A-F Experiments on Different Pre-trained Feature Extractor
To compute the cosine similarity in the proposed PFS, a pre-trained feature extractor $E$ is required. This section examines the impact of various pre-trained feature extractors $E$ on the CIFAR-10 dataset. As demonstrated in Figure 12, even using a common feature extractor pre-trained on the ImageNet dataset yields significant improvements compared to the baseline.
Figure 12: Experiments on Different Pre-trained Models. The attack success rate under different pre-trained models on the CIFAR-10 dataset. All results are computed the mean by 5 different run. Experiment settings are the same as that on the Table 1 of the original paper.

### v3

@@ A-E Attack Success Rate Under Additional Target Classes on the CIFAR-10 Dataset
A-F Experiments on CIFAR-100 Dataset
The results for CIFAR-100 in Table XII are similar to those for CIFAR-10. Our PFS approach outperforms random selection in all cases, with average boosts of 0.047, 0.068, and 0.037 for BadNets, Blended, and Optimized triggers, respectively. Compared to FUS, PFS achieves better attack results in 49 out of 54 settings.
TABLE XII: The attack success rate on CIFAR-100 dataset. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets, Blended, and Optimized are $1.5\%$, $1.5\%$, and $0.75\%$, respectively. Among four select strategies, the best result is denoted in boldface while the underline indicates the second-best result. The settings corresponding to the gray background are represented as the proxy poisoning attack within the attacker phase and the actual poisoning process within the victim phase being consistent in FUS.